OpenStudy (anonymous):

Consider integer solutions to the equation: \(x_1+x_2+x_3+x_4+x_5=10\) How many non-negative integer solutions are there such that \(x_1>x_2\)?

ganeshie8 (ganeshie8):

a dumb method wud be to consder cases x1 - x2 = 1,2,3,4,5,6,7,8,9

ganeshie8 (ganeshie8):

@Zarkon help

OpenStudy (zarkon):

use the previous answer and symmetry

OpenStudy (zarkon):

you need more hints?

OpenStudy (anonymous):

No. of non-negative integer solutions without restrictions = C(14,4) = 1001 No. of non-negative integer solutions with either \(x_1>x_2\) or \(x_2>x_1\) = C(14,4) - 161 = 1001-161 = 840 If you say use symmetry, I suppose no. of non-negative integer solutions with \(x_1>x_2\) = no. of non-negative integer solutions with \(x_1<x_2\)?! So, we get no. of non-negative integer solutions with \(x_1>x_2\) = 840/2 = 420? Ehhhh....

OpenStudy (zarkon):

correct

ganeshie8 (ganeshie8):

beautiful ! XD

OpenStudy (anonymous):

Thanks, but how can you explain that symmetry concept?

OpenStudy (zarkon):

you already did. there are 3 possibilities \(x_1>x_2,x_1=x_2,x_1<x_2\) by the symmetry of the problem \(x_1>x_2x_1<x_2\) have to have the same number of possibilities (you could switch the roles of \(x_1\) and \(x_2\) and you would still have the same number f solutions)

OpenStudy (zarkon):

or you could do it the long way \[\sum_{i=0}^{10}\sum_{j=i+1}^{10-i}{2+10-i-j\choose 2}=420\] ;)

OpenStudy (zarkon):

typo \[\sum_{i=0}^{4}\sum_{j=i+1}^{10-i}{2+10-i-j\choose 2}=420\]

OpenStudy (zarkon):

that would be going through the cases

OpenStudy (anonymous):

@RolyPoly , how would the total number of non-negative integer solutions without restriction be C(14,4)? I understand that C(14,4) means 14 choose 4, but I am curious where that came from.

OpenStudy (anonymous):

Stars and bars :D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Okay, give me a moment to chew on that. I am familiar with stars and bars.

OpenStudy (anonymous):

@ganeshie8 , thanks for the link -- I am enjoying reading!

OpenStudy (anonymous):

So, for the symmetry, after eliminating the solutions with \(x_1 = x_2\), we only have \(x_1>x_2\) or \(x_2>x_1\). Since we can switch the role of \(x_1\) and \(x_2\) without affecting the values of other variables, the number of solutions with \(x_1>x_2\) and with \(x_2>x_1\) are the same. Does this explanation looks good? Or still missing something?

OpenStudy (zarkon):

that looks fine

OpenStudy (zarkon):

you can think of it like this too.. you have x1 and x2...but I called them x2 and x1 (backwards from you) so even though we are using a different order ...we are really doing the same problem.

OpenStudy (anonymous):

Got it! Thanks!! :D